OpenStudy (anonymous):

Given that A is an angle and the answer to sec A=2.4 is sec(-1)(2.4), could it be converted to Cosine by secant A=1/cos A, multiply by 1 ie cos A/1 divided by cos A/1therefore secant (-1) = cos A??

OpenStudy (amistre64):

sec and cos are reciprocals of each other

OpenStudy (amistre64):

\[\frac{cos}{1}=\frac{1}{sec}\] \[\frac{sec}{1}=\frac{1}{cos}\]

OpenStudy (amistre64):

\[sec(A) = 2.4 \ \implies \ sec^{-1}(2.4) = A\] i cant say for sure, but I dont think the '-1' means flip it; basically becasue sec is the name of a function. f(x) = 3x+4 but does f^(-1)(x) = 1/(3x+4) ??

OpenStudy (amistre64):

to invert is to undo. the inverse of a function coupled with the function produce an identity

OpenStudy (amistre64):

\[(f\ o\ f^{-1})(x)=x\] if im remembering this right